OpenStudy (anonymous):

Use Cramer's rule to solve the system of equations. { 2x+3y=4 x-2y=9

OpenStudy (anonymous):

Are you studying matrices / linear algebra?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

Great, so how far have you got until now?

OpenStudy (anonymous):

(If you're stuck from the beginning, that's fine too though!)

OpenStudy (anonymous):

i understand how to do it, just the last step is where i mess up. I only know how to put them into the matrix sets, and solve those out. but i have no idea where to go from there.

OpenStudy (anonymous):

As I remember it, you create a square matrix A (left hand side of the equation) and column (that's vertical) vector b (right side of the equation). To get x, you replace the first column of your square matrix with vector b, whereas for y you replace the second column of your square matrix with vector b instead. In both cases, you then take the determinant of the new matrix you've created, and divide by the determinant of the original square matrix A. Does that make sense?

OpenStudy (anonymous):

If so, we can try it with the numbers now?

OpenStudy (anonymous):

If that doesn't make sense, we can do it using Draw.

OpenStudy (anonymous):

\[\left[\begin{matrix}2 & 3 \\ 1 & -2\end{matrix}\right]\left(\begin{matrix}x \\ y\end{matrix}\right)=\left(\begin{matrix}4 \\ 9\end{matrix}\right)\]

OpenStudy (anonymous):

yea i understand,

OpenStudy (anonymous):

Thanks iven5880!

OpenStudy (anonymous):

I'm still working on it. I'm not goot at using this editor

OpenStudy (anonymous):

So the original square matrix is the one that iven5880 has laid out, and whose determinant is 2*-2 -3*1 = -4-3=-7.

OpenStudy (anonymous):